OpenStudy (anonymous):

graph the hyperbola with equation ((x+2)^2/49)-((y+3)^2/4=1 PLEASE HELP WILL GIVE MEDALS

OpenStudy (anonymous):

\[((x+2)^2/49)-((y+3)^2/4)=1\]

OpenStudy (michele_laino):

we have to perform this traslation first: \[\left\{ \begin{gathered} X = x + 2 \hfill \\ Y = y + 3 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

where X, Y are the new coordinates

OpenStudy (michele_laino):

using those new coordinates, we can write your hyperbola as follows: \[\frac{{{X^2}}}{{{7^2}}} - \frac{{{Y^2}}}{{{2^2}}} = 1\]

OpenStudy (anonymous):

how do I graph it? I have a TI-84 and I cant figure it out

OpenStudy (michele_laino):

here is the graph of both coordinate systems: 

OpenStudy (anonymous):

wont they be parabolas?

OpenStudy (michele_laino):

in the XOY system the graph of your hyperbola is: